OpenStudy (anonymous):

One end of a rope is vibrated to produce a wavelength of 2.5 meters. The frequency of the wave is 5.0 Hz. What is the speed of the wave?

OpenStudy (anonymous):

\[v _{p}=\lambda/\tau\], \[v _{p}\] speed, \[\lambda\]=2.5m, \[\tau=1/f=1/5\], \[v _{p}\]=2.5/(1/5)=12,5m/s